# On quivers satisfying Quantum Yang-Baxter equation, and Hecke condition

### v1

@@ Abstract.
In this paper we initiate the study of quivers satisfying quantum Yang-Baxter equation. By considering quivers with quantum group like relations such as quantum Yang-Baxter equation, Hecke condition, we follow RTT construction to produce examples of Leavitt path algebras that contain quantum matrix algebra as subalgebra. Furthermore, We show that the quantum Yang-Baxter equation, and Hecke condition for our RTT construction can be generated intrinsically from the adjacency matrix of certain quivers.
@@ 1. Introduction
In Section 2 of this paper we introduce the notion of quiver satisfying the quantum Yang-Baxter equation. Our approach is quite different from the one in [2]. We say that a quiver $Q$ satisfies the QYBE if the adjacency matrix of the Kronecker square quiver satisfies the QYBE. We show that if a quiver $Q$ satisfies the QYBE then its adjacency matrix $A$ satisfies the “Temperley-Lieb idempotent” property, that is, $A^{2}=\mu A$ for some scalar $\mu$. In particular, we show that a finite symmetric quiver $Q$ satisfies the QYBE if and only if each nontrivial connected component of $Q$ is a complete weighted quiver (that is, there is a nonzero weighted arrow between any two vertices in the component).
@@
Closely linked to the QYBE is the Hecke condition, which places an additional polynomial constraint on the $R$-matrix. In its common form, the Hecke condition states that $(R-qI)(R+q^{-1}I)=0$ for some nonzero parameter $q$. Together, the QYBE and the Hecke condition provide a powerful algebraic framework used to construct quantum symmetries and invariants. Solutions that satisfy both conditions lead to braided monoidal categories, quantum knot invariants like the Jones polynomial, and integrable lattice models in statistical mechanics. In noncommutative algebra, these equations guide the formation of deformed algebras where commutation rules are governed by the $R$-matrix. In Section 3 of this paper, we give a recipe to construct Hecke $R$-matrices from the weighted adjacency matrices of certain quivers.
@@
Recently, there has been a lot of interest in quantum graphs among researchers working in quantum communication theory. Quantum graphs were first introduced in 1930s by Pauling to describe organic molecules and since then they have been studied from different perspectives. Physicists define it as a metric graph with the Schrödinger operator defined at the edges (according to the physical motivation) and some coupling conditions at the graph vertices that specify the transition probabilities across the vertices (see for example [8]). Weaver [13] proposed quantum graphs as reflexive symmetric quantum relations on a von Neumann algebra. In [4], quantum graphs are defined in terms of a self-adjoint operator on $L^{2}(H)$, for a finite-dimensional $C^{*}$-algebra $H$, which is idempotent under pointwise multiplication generalizing the usual adjacency matrix of a graph. Brannan at al adapt this notion in [5] to directed graphs and develop the notion of quantum graph $C^{*}$-algebras. Leavitt path algebras are algebraic analogues of graph C*-algebras defined as certain quotients of path algebras where the relations are inspired from Cuntz-Krieger relations for graph $C^{*}$-algebras [1]. Leavitt path algebras have deep connections with symbolic dynamics and since they capture the homological properties of both path algebras and their Koszul dual, they form an important class of noncommutative algebras. It is known that the Leavitt path algebra over Dynkin quiver $A_{n}$ is the matrix algebra $M_{n}(\mathbb{K})$ and so it is natural to ask if quantum matrix algebra can be realized as Leavitt path algebra over some quiver.
@@
In Section 4, we consider quivers with RTT relations determined by Hecke R-matrices and quantum Yang-Baxter equation and obtain quantum matrix algebra as subalgebra of the Leavitt path algebras over quivers with such relations. There is a natural operator-algebraic interpretation of the RTT-enhanced Leavitt path algebras constructed in this section. The imposition of local RTT relations at vertices with loop spaces replaces the free $C^{*}$-correspondence of [6] by a braided tensor correspondence determined by the Hecke $R$-matrix $R_{u}$. In the classical setting ($R_{u}=\mathrm{Flip}$), the correspondence is symmetric and yields the usual graph $C^{*}$-algebra. When $R_{u}$ is nontrivial, the tensor product over loop space $L(u)$ becomes a braided tensor product, placing $L_{\mathbb{K}}(Q,R)$ into the framework of Hayashi’s face-type quantum group correspondences [10] and more general quantum graph $C^{*}$-algebras. In particular, for a vertex $u$ with $\dim_{\mathbb{K}}L(u)=n^{2}$ and Hecke $R$-matrix arising from $GL_{q}(n)$, the subalgebra generated by the loop edges yields the quantum matrix algebra $\mathcal{O}_{q}(M_{n})$, and the resulting $L_{\mathbb{K}}(Q,R)$ may be viewed as a quantum graph algebra whose quantum symmetries are encoded combinatorially by the quiver and the RTT relations.
@@
Thus, the construction developed here produces natural algebraic models for quantum Cuntz–Krieger algebras and braided graph correspondences, enriching the interface between Leavitt path algebras, quantum groups, and noncommutative topology.

### v2

# Quantum spheres as Leavitt path algebras: Quivers with Quantum Yang-Baxter equation, and Hecke condition

@@ Abstract.
In this paper we study Leavitt path algebras over quivers with relations such as quantum Yang-Baxter equation, Hecke condition, and RTT conditions. This construction allows us to produce examples of Leavitt path algebras that contain quantum matrix algebra as subalgebra and obtain an algebraic analogue of the Hong-Szymański result. In particular, we show that the coordinate algebra over odd-dimensional Vaksman-Soibelman quantum sphere can be realized as the Zhang twist of a Leavitt path algebra over a quiver with such relations. Furthermore, We show that the quantum Yang-Baxter equation, and Hecke condition for our RTT construction can be generated intrinsically from the adjacency matrix of certain quivers.
Mathematics Subject Classifications 2020: 81R50, 16S88.
Key words: Quantum Yang-Baxter equation, Hecke Condition, Leavitt path algebra, Quantum sphere, Zhang twist.
@@ 1. Introduction
The central idea in noncommutative geometry is to characterize quantum spaces by their noncommutative algebra of functions. The first examples of quantum spaces are due to Podleś who introduced his well-known examples of 2-dimensional quantum spheres as homogeneous spaces for Woronowicz’s quantum $SU(2)$ group. The Vaksman-Soibelman quantum sphere, denoted as $S_{q}^{2n+1}$, represents a fundamental construction in the field of noncommutative geometry. These were introduced by Vaksman and Soibelman in the 1990s as quantum homogeneous spaces of higher dimensional quantum unitary groups. For $n\in\mathbb{N}$ and $q\in[0,1]$, a deformation parameter, the $(2n+1)$-dimensional quantum sphere $S_{q}^{2n+1}$ is described by a filtered quadratic $\mathbb{C}$-algebra. For $n=0$, one gets the algebra of polynomial functions on the unit circle, and for $n=1$ the quantum space underlying Woronowocz’s quantum unitary group $SU(2)$. Hong and Szymański [13] proved that the enveloping $C^{*}$-algebra $C(S_{q}^{2n+1})$ of the odd-dimensional Vaksman-Soibelman quantum sphere $S_{q}^{2n+1}$ is isomorphic to the graph $C^{*}$-algebra $C^{*}(L_{2n+1})$ over a quiver $L_{2n+1}$. See D́ Andrea’s article [3] for a detailed survey of this topic. Leavitt path algebras are algebraic analogues of graph $C^{*}$-algebras defined as certain quotients of path algebras where the relations are inspired from Cuntz-Krieger relations for graph $C^{*}$-algebras [1]. Leavitt path algebras have deep connections with symbolic dynamics and since they capture the homological properties of both path algebras and their Koszul dual, they form an important class of noncommutative algebras. The primary objective of this paper is to find an algebraic analogue of the Hong-Szymański result, and realize the coordinate algebra over odd-dimensional Vaksman-Soibelman quantum sphere as the Leavitt path algebra over some quiver. It turns out that in this context we need to consider quivers with some quantum group like relations such as Quantum Yang-Baxter Equation, Hecke condition and RTT condition.
@@
In Section 2 of this paper we extend the study of QYBE to another category by introducing the concept of quivers satisfying the Quantum Yang-Baxter Equation. Our approach is quite different from the one in [3]. We say that a quiver $Q$ satisfies the QYBE if the adjacency matrix of the Kronecker square quiver satisfies the QYBE. We show that if a quiver $Q$ satisfies the QYBE then its adjacency matrix $A$ satisfies the “Temperley-Lieb idempotent” property, that is, $A^{2}=\mu A$ for some scalar $\mu$. In particular, we show that a finite symmetric quiver $Q$ satisfies the QYBE if and only if each nontrivial connected component of $Q$ is a complete weighted quiver (that is, there is a nonzero weighted arrow between any two vertices in the component).
@@
Closely linked to the QYBE is the Hecke condition, which places an additional polynomial constraint on the $R$-matrix. The Hecke condition states that $(R-qI)(R+q^{-1}I)=0$ for some nonzero parameter $q$. Together, the QYBE and the Hecke condition provide a powerful algebraic framework used to construct quantum symmetries and invariants. Solutions that satisfy both conditions lead to braided monoidal categories, quantum knot invariants like the Jones polynomial, and integrable lattice models in statistical mechanics. In noncommutative algebra, these equations guide the formation of deformed algebras where commutation rules are governed by the $R$-matrix. In Section 3 of this paper, we give a recipe to construct Hecke $R$-matrices from the weighted adjacency matrices of certain quivers.
@@
In Section 4, we consider quivers with RTT relations determined by Hecke R-matrices and quantum Yang-Baxter equation and study Leavitt path algebras over quivers with such relations. Recently, there has been a lot of interest in quantum graphs among researchers working in quantum communication theory. Quantum graphs were first introduced in 1930s by Pauling to describe organic molecules and since then they have been studied from different perspectives. Physicists define it as a metric graph with the Schrödinger operator defined at the edges (according to the physical motivation) and some coupling conditions at the graph vertices that specify the transition probabilities across the vertices (see for example [10]). Weaver [18] proposed quantum graphs as reflexive symmetric quantum relations on a von Neumann algebra. In [6], quantum graphs are defined in terms of a self-adjoint operator on $L^{2}(H)$, for a finite-dimensional $C^{*}$-algebra $H$, which is idempotent under pointwise multiplication generalizing the usual adjacency matrix of a graph. Brannan at al adapt this notion in [7] to directed graphs and develop the notion of quantum graph $C^{*}$-algebras. Since Leavitt path algebras are algebraic counterparts of graph $C^{*}$-algebras, it is natural to ask if we can obtain quantum matrix algebras as examples of our construction here. In Example 4.3 we obtain quantum matrix algebra $\mathcal{O}_{q}(M_{n})$ as a subalgebra of a Leavitt path algebra over a quiver with such relations.
@@
There is a natural operator-algebraic interpretation of the Leavitt path algebras with RTT relations constructed in this section. The imposition of local RTT relations at vertices with loop spaces replaces the free $C^{*}$-correspondence of [8] by a braided tensor correspondence determined by the Hecke $R$-matrix. In the classical setting, the correspondence is symmetric and yields the usual graph $C^{*}$-algebra. When the $R$-matrix is nontrivial, the tensor product over loop space becomes a braided tensor product, placing $L_{\mathbb{K}}(Q,R)$ into the framework of quantum graph $C^{*}$-algebras. In particular, for a vertex $u$ with the loop space $L(u)$ having dimension $n^{2}$ and Hecke $R$-matrix arising from $GL_{q}(n)$, the subalgebra generated by the loop edges yields the quantum matrix algebra $\mathcal{O}_{q}(M_{n})$, and the resulting $L_{\mathbb{K}}(Q,R)$ may be viewed as a quantum graph algebra whose quantum symmetries are encoded combinatorially by the quiver and the RTT relations.
In Section 5, we prove our main result, an algebraic analogue of the Hong-Szymański result, by showing that the coordinate algebra over odd-dimensional Vaksman-Soibelman quantum sphere can be realized as the Zhang twist of a Leavitt path algebra over quiver $L_{2n+1}$ with QYBE, Hecke condition and RTT relations. Recall that a notion of twist of a graded algebra $A$ was introduced by Artin, Tate, and Van den Bergh in [4] as a deformation of the original graded product of $A$ with the help of a graded automorphism of $A$. Let $\sigma$ be an automorphism of the graded algebra $A=\oplus A_{n}$. Define a new multiplication $\star$ on the underlying graded $K$-module $\oplus A_{n}$ by $a\star b=a\sigma^{n}(b)$ where $a$ and $b$ are homogeneous elements in $A=\oplus A_{n}$ and $deg(a)=n$. The new graded algebra with the same underlying graded $K$-module $\oplus A_{n}$ and the new graded product $\star$ is called the twist of $A$ and is denoted as $A^{\sigma}$. This notion of twist of a graded algebra was later generalized by Zhang in [21], where he introduced the concept of twisting of graded product with the help of a twisting system. We refer the readers to [15] for more details on Zhang twist of Leavitt path algebras and their applications.
@@
Thus, the construction developed here produces natural algebraic models for quantum Cuntz–Krieger algebras and braided graph correspondences, enriching the interface between Leavitt path algebras, quantum groups, and noncommutative geometry.
@@ 4. Application: Leavitt path algebra over quiver with relations
5. Vaksman-Soibelman Quantum Sphere as Leavitt Path Algebra over quivers with relations
It is known due to Hong and Szymański [13] that $C^{*}$-algebra $C(S_{q}^{2n+1})$ of the odd-dimensional Vaksman-Soibelman quantum sphere $S_{q}^{2n+1}$ is isomorphic to the graph $C^{*}$-algebra $C^{*}(L_{2n+1})$ where $L_{2n+1}$ is the quiver with vertex set
$\{v_{0},v_{1},\dots,v_{n}\}$
and arrow set
$\{e_{ij}:v_{i}\to v_{j}\mid 0\leq i\leq j\leq n\}.$
We proceed to find an algebraic analogue of this result. Let $K$ be a field and let $q\in K^{\times}$. Fix an integer $n\geq 1$. Then $\mathcal{O}(S_{q}^{2n+1})$ is defined as the associative $K$-algebra generated by $Z_{0},Z_{1},\ldots,Z_{n}$, $Z_{0}^{*},Z_{1}^{*},\ldots,Z_{n}^{*}$ with relations
$Z_{j}Z_{i}=qZ_{i}Z_{j}$ for $0\leq i<j\leq n$
$Z_{i}^{*}Z_{j}=qZ_{j}Z_{i}^{*}$ for $0\leq i\neq j\leq n$
$Z_{i}^{*}Z_{i}=Z_{i}Z_{i}^{*}+(1-q^{2})\sum_{j=i+1}^{n}Z_{j}Z_{j}^{*}$ for $0\leq i\leq n$.
Let $Q=L_{2n+1}$ be the quiver with vertex set
$Q_{0}=\{v_{0},v_{1},\dots,v_{n}\}$
and arrow set
$Q_{1}=\{e_{ij}:v_{i}\to v_{j}\mid 0\leq i\leq j\leq n\}.$
Let $\overline{Q}$ be the double quiver with ghost arrows $e_{ij}^{*}$.
Let $E(v_{0}):=\mathrm{span}_{K}\{e_{0j}\mid 0\leq j\leq n\}$ and let $F:E(v_{0})\otimes E(v_{0})\to E(v_{0})\otimes E(v_{0})$ be the flip map $F(x\otimes y)=y\otimes x$. Define
$R_{v_{0}}:=F\in\operatorname{End}(E(v_{0})\otimes E(v_{0})).$
Then $R_{v_{0}}$ satisfies QYBE and the Hecke relation with parameter $1$: $(R_{v_{0}}-I)(R_{v_{0}}+I)=0$.
Impose the RTT relations at $v_{0}$.
(10) $e_{0j}e_{0i}=e_{0i}e_{0j},\qquad e_{0i}e_{0j}^{*}=e_{0j}^{*}e_{0i},\qquad 0\leq i<j\leq n,$
Let $L$ be the Leavitt path algebra over quiver $Q=L_{2n+1}$ with RTT relations. Let $p:=v_{0}\in L$ and consider the corner $pLp$.
Define a graded automorphism $\sigma:L\to L$ by
$\sigma(v_{i})=v_{i},\qquad\sigma(e_{ii})=e_{ii}\ (0\leq i\leq n),\qquad\sigma(e_{0j})=q\,e_{0j}\ (1\leq j\leq n),\qquad\sigma(e_{0j}^{*})=q^{-1}e_{0j}^{*},$
and $\sigma(e_{ij})=e_{ij}$, $\sigma(e_{ij}^{*})=e_{ij}^{*}$ for all remaining arrows with $1\leq i<j\leq n$.
We know that $L$ has a $\mathbb{Z}$-graded structure as $L=\oplus_{n}L_{n}$. We twist the multiplicative structure of $\oplus_{n}L_{n}$ as $a\star b=a\sigma^{n}(b)$ for any $a\in L_{n},b\in L_{m}$. The same underlying graded vector space $\oplus L_{n}$ with this new graded product $\star$ is called the Zhang twist of $L$ and denoted as $L^{\sigma}$.
We write $pL^{\sigma}p$ for the twisted corner algebra. Define elements in $pL^{\sigma}p$ by
$z_{0}:=e_{00},\qquad z_{k}:=e_{0k}e_{kk}e_{0k}^{*}\quad(1\leq k\leq n).$
Let $J_{q}$ be the ideal of $pL^{\sigma}p$ generated by the relations
(11) $z_{i}^{*}\star z_{i}=z_{i}\star z_{i}^{*}+(1-q^{2})\sum_{k=i+1}^{n}z_{k}\star z_{k}^{*},\qquad i=0,1,\dots,n-1.$
Theorem 5.1.
The quotient algebra
$\mathcal{A}_{q}^{2n+1}:=\big(pL^{\sigma}p\big)/J_{q}$
is isomorphic to the algebra $\mathcal{O}(S_{q}^{2n+1})$ of the odd-dimensional Vaksman–Soibelman quantum sphere. More precisely, the assignment
$\Phi:\mathcal{O}(S_{q}^{2n+1})\to\mathcal{A}_{q}^{2n+1},\qquad\Phi(Z_{i})=\overline{z_{i}},$
where $\{Z_{0},\dots,Z_{n}\}$ are the standard generators of $\mathcal{O}(S_{q}^{2n+1})$ and $\overline{z_{i}}$ are the images in the quotient, is a well-defined isomorphism.
Proof.
We prove that the elements $z_{0},\dots,z_{n}$ satisfy the defining relations of $\mathcal{O}(S_{q}^{2n+1})$ in the quotient $\mathcal{A}_{q}^{2n+1}$, and then show the universal map is an isomorphism.
By definition $R_{v_{0}}=F$ is the flip on $E(v_{0})\otimes E(v_{0})$. The flip is a standard solution of QYBE, and it satisfies the Hecke relation with parameter $1$:
(FI)(F+I)=0(F-I)(F+I)=0
since $F^{2}=I$.
We work in the untwisted corner $pLp$ first. For each $k\geq 1$,
$z_{k}z_{k}^{*}=(e_{0k}e_{kk}e_{0k}^{*})(e_{0k}e_{kk}^{*}e_{0k}^{*})=e_{0k}(e_{kk}e_{kk}^{*})e_{0k}^{*}.$
At the vertex $v_{k}$, the only outgoing edge is the loop $e_{kk}$, hence CK2 gives $e_{kk}e_{kk}^{*}=v_{k}$ and so
$z_{k}z_{k}^{*}=e_{0k}e_{0k}^{*}.$
Also $z_{0}z_{0}^{*}=e_{00}e_{00}^{*}$. At $v_{0}$ the outgoing edges are $e_{00},e_{01},\dots,e_{0n}$, hence CK2 yields
$p=v_{0}=e_{00}e_{00}^{*}+\sum_{k=1}^{n}e_{0k}e_{0k}^{*}.$
Substituting $e_{0k}e_{0k}^{*}=z_{k}z_{k}^{*}$ gives
(12) $z_{0}z_{0}^{*}+\sum_{k=1}^{n}z_{k}z_{k}^{*}=p.$
Because $\sigma(z_{i})=z_{i}$ for all $i$ (the $q$-factors cancel in $e_{0k}(\cdot)e_{0k}^{*}$, and $e_{00}$ is fixed), the Zhang twist does not change (12); hence in the twisted corner we still have
$z_{0}\star z_{0}^{*}+\sum_{k=1}^{n}z_{k}\star z_{k}^{*}=p.$
In $L$, the RTT relations (10) imply in particular
$e_{0j}e_{0i}=e_{0i}e_{0j}\qquad(i<j).$
We compute in the twist $L^{\sigma}$: for $i<j$ with $1\leq i,j\leq n$,
$e_{0j}\star e_{0i}=e_{0j}\,\sigma(e_{0i})=e_{0j}\,(qe_{0i})=q\,e_{0j}e_{0i}=q\,e_{0i}e_{0j}=e_{0i}\,\sigma(e_{0j})=e_{0i}\star e_{0j}.$
Thus
(13) $e_{0j}\star e_{0i}=e_{0i}\star e_{0j}\qquad(1\leq i<j\leq n),$
and similarly, using the RTT commutation $e_{0i}e_{0j}^{*}=e_{0j}^{*}e_{0i}$ and $\sigma(e_{0j}^{*})=q^{-1}e_{0j}^{*}$,
$e_{0i}\star e_{0j}^{*}=e_{0i}\sigma(e_{0j}^{*})=e_{0i}(q^{-1}e_{0j}^{*})=q^{-1}e_{0i}e_{0j}^{*}=q^{-1}e_{0j}^{*}e_{0i}=e_{0j}^{*}\sigma(e_{0i})=e_{0j}^{*}\star e_{0i}.$
These are precisely the $q$-commutation relations once we pass from the edge generators to the corner generators $z_{i}$ defined by $z_{i}=e_{0i}e_{ii}e_{0i}^{*}$: the relations (13) and their ghost analogues imply the standard Vaksman–Soibelman commutation relations among the $z_{i}$.
The remaining defining relations of $\mathcal{O}(S_{q}^{2n+1})$ are given by relations
$Z_{i}^{*}Z_{i}=Z_{i}Z_{i}^{*}+(1-q^{2})\sum_{k=i+1}^{n}Z_{k}Z_{k}^{*}\qquad(i=0,\dots,n-1).$
By construction, in $\mathcal{A}_{q}^{2n+1}=(pL^{\sigma}p)/J_{q}$ the elements $\overline{z_{i}}$ satisfy these relations because $J_{q}$ is precisely the ideal generated by (11). This shows that the assignment $Z_{i}\mapsto\overline{z_{i}}$ respects all defining relations of $\mathcal{O}(S_{q}^{2n+1})$. Hence it extends to a well-defined surjective homomorphism
$\Phi:\mathcal{O}(S_{q}^{2n+1})\twoheadrightarrow\mathcal{A}_{q}^{2n+1}.$
Conversely, $\mathcal{A}_{q}^{2n+1}$ is generated by $\overline{z_{0}},\dots,\overline{z_{n}}$, and the only relations imposed among these generators are exactly the sphere relation, the $q$-commutation relations, and the relations (11). Thus $\mathcal{A}_{q}^{2n+1}$ has the same presentation as $\mathcal{O}(S_{q}^{2n+1})$. By the universal property of presented algebras, there exists a homomorphism
$\Psi:\mathcal{A}_{q}^{2n+1}\to\mathcal{O}(S_{q}^{2n+1})$
sending $\overline{z_{i}}\mapsto Z_{i}$. The compositions $\Psi\circ\Phi$ and $\Phi\circ\Psi$ act as the identity on the respective generating sets, hence are identity maps. Therefore $\Phi$ is an isomorphism. ∎